OpenStudy (anonymous):

Find the sum of all real solutions of the equation x^2−16=3(x−4)^2.

OpenStudy (anonymous):

\[x^{2}-16=3(x-4)^2\] using difference of two squares we simplify left hand side \[x^2-4^2=(x-4)(x+4)\] \[(x-4)(x+4)=3(x-4)(x-4)\] \[x+4=3(x-4)\]

OpenStudy (anonymous):

I know that . But I got the Answer x=8 and its wrong .

OpenStudy (anonymous):

that implies the above method is wrong,let me think of another way

OpenStudy (anonymous):

Ok thanks , I'll wait :